OpenStudy (anonymous):

hi, so the directions are to solve the exponential expression without using a calculator... my question is 2^3x-2=128. how do i even start this?!?!??!

OpenStudy (anonymous):

is this the problem? \[2^{3x}-2=128\]

OpenStudy (anonymous):

\[2^{3x}-2=128\] This can be re-written as: \[2^{3x}-2^1=2^7\] 3x-1=7 Solve for x x=8/3